Please answer step by step: Two firms compete in a market to sell a homogeneous product with inverse demand function P= 400-2Q. Each firm produces at a constant marginal cost of $50 and has no fixed costs. Use this information to compare the output levels and profits in settings characterized by Cournot, Stackelberg, Bertrand, and collusive behavior.
Answer #1

- Bertrand P= 400-26 MC = 50 P=mc Condition under the bestrand competition. 400-28 = 50 T P = [50] 350 = 20 8 = 350 I r= 10 2Monopoly P= 400-2ą TR = 4008-20 MR= 400-48 P = 400-2X 87.5 = 400 175 = 225 c = so MR= MC 400-48=50 300 = 48 &= 350 4 = 87.5 r